OpenStudy (anonymous):

help out x-3x=-13 4x+2y=4 what's the (x,y)?

OpenStudy (a1234):

There are three ways you can solve systems of equations: graphing, elimination, and substitution. Do you know how to use these methods?

OpenStudy (anonymous):

1st problem , X=13/2 2nd problem , X = 1-1/2y

OpenStudy (a1234):

Do you have a typo in the first equation?

OpenStudy (anonymous):

thankz and okay and no i dont have a typo in the first question thats how it is

OpenStudy (anonymous):

@a1234 which method do you highly suggest i should use?

OpenStudy (a1234):

Substitution.

OpenStudy (anonymous):

okay thanks

OpenStudy (a1234):

Sure.

OpenStudy (anonymous):

:)